OpenStudy (anonymous):

sqrt 3 y^2 + 4 sqrt 12 y^2 - y sqrt 75

OpenStudy (gorv):

\[\sqrt{3*y^2}+4\sqrt{12*y^2}-y*\sqrt{75}\]

OpenStudy (gorv):

\[\sqrt{3}*y+4\sqrt{3*4}*y-y*\sqrt{25*3}\]

OpenStudy (gorv):

can u solve it now??

OpenStudy (anonymous):